# A several variables Kowalski-Słodkowski theorem for topological spaces

### v1

@@ 2. $(\mathcal{W}2)$
[10, Theorem 1.4] Let $\mathcal{X}$ be a Banach space that satisfies the above properties $(JS1)-(JS3)$ over an open set $U\subset\mathbb{C}^{n}$. Let $\Lambda:\mathcal{X}\to\mathbb{C}$ be a continuous linear functional such that $\Lambda(e^{w\cdot z})\neq 0$ for every $w\in\mathbb{C}^{n}$. Then there exist a non-zero $\alpha\in\mathbb{C}$ and $w\in U$ such that
@@
We now focus on an analogue of the KS theorem for function spaces. By following the ideas of Mashreghi and Ransford from [8], Sebastian and Daniel extended the KS theorem to modules in [11, Theorem 2.3]; then they used it to obtain the following generalization of Theorem 1.3.
@@
[11, Theorem 3.1] Let $1\leq p\leq\infty$ and $F:H^{p}(\mathbb{D})\rightarrow\mathbb{C}$ be a function such that $F(0)=0,\ F(1)\neq 0,$ and
@@
    Theorem 1.6 ([10, Theorem 1.4]) is a several variables version of the GKZ theorem, whereas our result (Theorem 3.2) is a several variables version of the KS theorem. In addition, our functions are vector-valued as well.
@@ 2. Kowalski-Słodkowski’s Theorem for Hardy spaces
The proof of Theorem 1.7 as given in [11] uses [11, Theorem 2.3]. We note that even for $0<p<1,$ the triple, $\mathcal{A}=H^{\infty}(\mathbb{D}),\ \mathcal{M}=H^{p}(\mathbb{D}),\ S$ equals the set of all outer functions in $H^{p}(\mathbb{D})$ satisfy the hypotheses of Theorem 2.3 from [11], and so just as shown in [11], there exists a character $\chi:H^{\infty}(\mathbb{D})\to\mathbb{C}$ such that $F(fg)=\chi(f)F(g)$ for all $f\in H^{\infty}(\mathbb{D})$ and $g\in H^{p}(\mathbb{D})$, where $0<p\leq\infty$. Further, by using arguments as were used in Theorem 1.7, there exists a point $w\in\mathbb{D}$ such that
@@
For each fixed $w\in\mathbb{D}$, the map $T_{w}:H^{p}(\mathbb{D})\to\mathbb{C}$ given by $T_{w}(f)=(Tf)(w)$ satisfies the hypotheses of Theorem 2.1; therefore, there exist a point $a_{w}\in\mathbb{D}$ and a constant $c_{w}$ such that $(Tf)(w)=c_{w}f(a_{w})$ for every $f\in H^{p}(\mathbb{D}).$ Then by taking $f$ to be the constant function $1$, we see that $c_{w}=T1(w)\neq 0.$ Further, by taking $f(z)=z$, we see $a_{w}=(Tz)(w)/T1(w).$ We now define $\psi=T1$ and $\phi=(Tz)/\psi$. Then $\psi$ and $\phi$ are holomorphic function on $\mathbb{D}$ such that $\psi$ doesn’t vanish at any point in $\mathbb{D}$, $\phi(\mathbb{D})\subseteq\mathbb{D}$, and $Tf=\psi\cdot(f\circ\phi)$. This completes the proof.
@@ 3. Kowalski-Słodkowski’s theorem for topological spaces
where $|\alpha|=\alpha_{1}+\alpha_{2}+\dotsm+\alpha_{n}$. For convenience, we shall write $\mathcal{P}_{n}$ instead of $\mathcal{P}_{n}^{\mathcal{A}}$ if $\mathcal{A}=\mathbb{C}.$ Clearly, given any compact subset $K$ of $\mathbb{C}^{n}$, $\mathcal{P}^{\mathcal{A}}_{n}$ can be viewed as a subalgebra of the Banach algebra $C(K,\mathcal{A})$, the set of all continuous $\mathcal{A}$-valued functions with the sup norm. We say $z_{0}$ is zero of a polynomial $p\in\mathcal{P}_{n}^{\mathcal{A}}$ if $p(z_{0})=0$. Given a set $D$ in $\mathbb{C}^{n}$, we use $\mathcal{P}_{n}^{0}(D)$ to denote the set of polynomials in $\mathcal{P}^{\mathcal{A}}_{n}$ with no zeros in $D$. For the rest of this section, let $\mathcal{A}$ represents a fixed complex Banach algebra with identity $e$.
The following result is integral to our version of the KS theorem for topological spaces (Theorem 3.2). We present it as an independent result as, apart from being interesting in its own right, it brings out the crux of the method used to prove Theorem 3.2 even more prominently. It characterizes multiplicative linear functionals on $\mathcal{P}^{n}_{\mathcal{A}}$ as non-zero complex-valued functions on $\mathcal{P}^{n}_{\mathcal{A}}$ that satisfy a condition very similar to the one assumed in the KS theorem. First, we recall the notion of a polynomial convex hull of a set in $\mathbb{C}^{n}.$ The polynomial convex hull $\widehat{D}$ of a bounded subset $D$ of $\mathbb{C}^{n}$ is consisting of all $w\in\mathbb{C}^{n}$ such that
@@
Let $\Lambda:\mathcal{P}^{n}_{\mathcal{A}}\to\mathbb{C}$ be a non-zero function with $\Lambda(0)=0$. Suppose $D$ is a bounded subset of $\mathbb{C}^{n}$ such that
$\big{(}\Lambda(p)-\Lambda(q)\big{)}e-(p-q)\notin\mathcal{P}^{0}_{n}(D)$
for all polynomials $p$ and $q$ in $\mathcal{P}^{\mathcal{A}}_{n}$. Then there exist $w\in\widehat{D}$ and a multiplicative linear functional $\phi$ on $\mathcal{A}$ such that
@@
Moreover, we can assert $w\in D$ whenever
@@
It is clear from the hypotheses that for every $p,\ q\in\mathcal{P}_{n}^{\mathcal{A}}$ there exists $w_{p,q}\in D$ such that $(\Lambda(p)-\Lambda(q))e=p(w_{p,q})-q(w_{p,q})$. In particular, for each polynomial $p$, there is a point $w_{p}$ in $D$ such that $\Lambda(p)e=p(w_{p})$. This gives us
@@
Now we want to show that $\big{(}\Lambda_{1}(f)-\Lambda_{1}(g)\big{)}e-(f-g)\notin G(\overline{\mathcal{P}_{n}^{\mathcal{A}}})$ (set of invertible elements of $\overline{\mathcal{P}_{n}^{\mathcal{A}}}$) for all $f,g\in\overline{\mathcal{P}_{n}^{\mathcal{A}}}$. Suppose there are $f$ and $g$ in $\overline{\mathcal{P}_{n}^{\mathcal{A}}}$ such that $\big{(}\Lambda_{1}(f)-\Lambda_{1}(g)\big{)}e-(f-g)\in G(\overline{\mathcal{P}_{n}^{\mathcal{A}}})$. Let $\{p_{n}\}_{n}$ and $\{q_{n}\}_{n}$ be sequences of polynomials in $\mathcal{P}_{n}^{\mathcal{A}}$ such that $p_{n}\longrightarrow f$ and $q_{n}\longrightarrow g$ in $\overline{\mathcal{P}_{n}^{\mathcal{A}}}$. Then
@@
Finally, we will establish the assertions made in the moreover part. First, suppose $n=1,$ that is, we are in the one-variable situation. Take $p(z)={\bf e}z,$ then $\Lambda(p)=w.$ Also, as we noted at the beginning of the proof, there exists a point $a\in D$ such that $\Lambda(p){\bf e}=p(a)={\bf e}a$, which implies that $w=a\in D.$
We now turn to the general $n$ case. Let $D=D_{1}\times\cdots\times D_{n}$ for some subsets $D_{1},\dots,D_{n}$ of $\mathbb{C}.$ The idea here is very similar to case $n=1$. Note that $\Lambda(r_{i})=\phi(r_{i}(w))=\phi(w_{i}e)=w_{i}$. Additionally, as mentioned earlier, some $a$ exists in $D$ such that $\Lambda(r_{i}){\bf e}=r_{i}(a)={\bf e}a_{i}.$ Thus $w_{i}=a_{i}\in D_{i}$; hence, $w\in D_{1}\times\cdots\times D_{n}=D.$
@@
Also, there exists some $a\in D$ such that $\Lambda(p){\bf e}=p(a)$, which implies that $||w-b||^{2}_{\mathbb{C}^{n}}=\sum_{i=1}^{n}(\overline{w_{i}-b_{i}})(a_{i}-b_{i})=\langle{w-b,a-b}\rangle_{\mathbb{C}^{n}}$. Thus, using the Cauchy-Schwarz inequality, we infer $||w-b||_{\mathbb{C}^{n}}\leq||a-b||_{\mathbb{C}^{n}}<r.$ Thus, $w\in D$. A similar set of arguments will work when $D$ is a closed Euclidean ball in $C^{n}$. This completes the proof. ∎
@@
Let $\Lambda:\mathcal{X}\to\mathbb{C}$ be a non-zero continuous function such that $\Lambda({0})=0$ and
$\big{(}\Lambda(p)-\Lambda(q)\big{)}e-(p-q)\notin\mathcal{P}^{0}_{n}(D)$
for every polynomial $p$ and $q$ in $\mathcal{P}_{n}^{\mathcal{A}}$. Then there exist $w\in D$ and a multiplicative linear functional $\phi$ on $\mathcal{A}$ such that
@@
Clearly, the restriction $\Lambda:\mathcal{P}_{n}^{\mathcal{A}}\longrightarrow\mathbb{C}$ satisfies the hypotheses of Theorem 3.1. Then, using the moreover part of Theorem 3.1, there exist $w\in D$ and a multiplicative linear functional $\phi$ on $\mathcal{A}$ such that $\Lambda(p)=\phi(p(w))$. Now the continuity of $\Lambda$ and the denseness of $\mathcal{P}_{n}^{\mathcal{A}}$ in $\mathcal{X}$ yield
@@
We would like to now compare Theorem 3.2 with Theorem 1.6 ([10, Theorem 1.4]) that also is a several variables analogue of the GKZ theorem for function spaces. First, we note that Theorem 1.6 is for Banach spaces of scalar-valued functions, whereas Theorem 3.2 deals with topological spaces of vector-valued functions. Next, we compare the domains of the function spaces considered by the two theorems. Theorem 3.2, in the several variables case, is proved under the assumption that the domain is either a Euclidean ball or a Cartesian product. In contrast, it may seem that the domain of functions in Theorem 1.6 includes a big class of subsets of $\mathbb{C}^{n}$. But, we want to point out that Theorem 1.6 assumes the domain to be a “maximal domain” (hypothesis JS2 in Theorem 1.6), and if we assume the domain $D$ in Theorem 3.2 to be a maximal domain instead of being a specific set such as a Euclidean ball or a Cartesian product (as done in Theorem 3.2), then indeed with the help of Theorem 3.1, we would also obtain the map $\Lambda$ to be a point evaluation for the scalar-valued case and composition of a multiplicative linear functional and a point evaluation for the vector-valued case. Moreover, Theorem 1.6 doesn’t cover Dirichlet spaces $\mathcal{D}_{\alpha}$ for $\alpha>1$ as the maximal domain for functions in these Dirichlet spaces is $\overline{\mathbb{D}^{n}}$ and not $\mathbb{D}^{n}$. But, these are covered by Theorem 3.2 as they satisfy its hypotheses.
@@
For $n,m\in\mathbb{N}$ and $0<p\leq\infty$, let $\Lambda:H^{p}(\mathbb{D}^{n},\mathbb{C}^{m})\to\mathbb{C}$ be a non-zero continuous (weak-star continuous for $p=\infty$) function such that $\Lambda({0})=0$ and $\big{(}\Lambda(p)-\Lambda(q)\big{)}{\bf 1}-(p-q)\notin\mathcal{P}^{0}_{n}(\mathbb{D}^{n})$ for every polynomial $p$ and $q$ in $\mathcal{P}^{\mathbb{C}^{m}}_{n}$. Then there exist $w\in\mathbb{D}^{n}$ and $1\leq k\leq m$ such that
@@
For $0<p,q\leq\infty,$ let $T:H^{p}(\mathbb{D}^{n})\to H^{q}(\mathbb{D}^{m})$ be a continuous function, where we consider weak-start topology on $H^{p}(\mathbb{D}^{n})$ when $p=\infty.$ Suppose $T(0)=0,\ T(1)(w)\neq 0$ for each $w\in\mathbb{D}^{m}$ and
$\big{(}Tp(w)-Tq(w)\big{)}1-\big{(}T1(w)\big{)}(p-q)\notin\mathcal{P}_{n}^{0}(\mathbb{D}^{n})$
for every $w\in\mathbb{D}^{m},$ and complex-valued polynomials $p,q$ in $n$-variables. Then there exist $\psi\in H^{q}(\mathbb{D}^{m})$ and an analytic function $\phi:\mathbb{D}^{m}\to\mathbb{D}^{n}$ such that

### v3

@@ 2. $(\mathcal{W}2)$
[11, Theorem 1.4] Let $\mathcal{X}$ be a Banach space that satisfies the above properties $(JS1)-(JS3)$ over an open set $U\subset\mathbb{C}^{n}$. Let $\Lambda:\mathcal{X}\to\mathbb{C}$ be a continuous linear functional such that $\Lambda(e^{w\cdot z})\neq 0$ for every $w\in\mathbb{C}^{n}$. Then there exist a non-zero $\alpha\in\mathbb{C}$ and $w\in U$ such that
@@
We now focus on an analogue of the KS theorem for function spaces. By following the ideas of Mashreghi and Ransford from [8], Sebastian and Daniel extended the KS theorem to modules in [12, Theorem 2.3]; then they used it to obtain the following generalization of Theorem 1.3.
@@
[12, Theorem 3.1] Let $1\leq p\leq\infty$ and $F:H^{p}(\mathbb{D})\rightarrow\mathbb{C}$ be a function such that $F(0)=0,\ F(1)\neq 0,$ and
@@
    Theorem 1.6 ([11, Theorem 1.4]) is a several variables version of the GKZ theorem, whereas our result (Theorem 3.2) is a several variables version of the KS theorem. In addition, our functions are vector-valued as well.
@@ 2. Kowalski-Słodkowski’s Theorem for Hardy spaces
The proof of Theorem 1.7 as given in [12] uses [12, Theorem 2.3]. We note that even for $0<p<1,$ the triple, $\mathcal{A}=H^{\infty}(\mathbb{D}),\ \mathcal{M}=H^{p}(\mathbb{D}),\ S$ equals the set of all outer functions in $H^{p}(\mathbb{D})$ satisfy the hypotheses of Theorem 2.3 from [12], and so just as shown in [12], there exists a character $\chi:H^{\infty}(\mathbb{D})\to\mathbb{C}$ such that $F(fg)=\chi(f)F(g)$ for all $f\in H^{\infty}(\mathbb{D})$ and $g\in H^{p}(\mathbb{D})$, where $0<p\leq\infty$. Further, by using arguments as were used in Theorem 1.7, there exists a point $w\in\mathbb{D}$ such that
@@
For each fixed $w\in\mathbb{D}$, the map $T_{w}:H^{p}(\mathbb{D})\to\mathbb{C}$ given by $T_{w}(f)=(Tf)(w)$ satisfies the hypotheses of Theorem 2.1; therefore, there exist a point $a_{w}\in\mathbb{D}$ and a constant $c_{w}$ such that $(Tf)(w)=c_{w}f(a_{w})$ for every $f\in H^{p}(\mathbb{D}).$ Then by taking $f$ to be the constant function $1$, we see that $c_{w}=T1(w)\neq 0.$ Further, by taking $f(z)=z$, we see $a_{w}=(Tz)(w)/T1(w).$ We now define $\psi=T1$ and $\phi=(Tz)/\psi$. Then $\psi$ and $\phi$ are holomorphic function on $\mathbb{D}$ such that $\psi$ doesn’t vanish at any point in $\mathbb{D}$, $\phi(\mathbb{D})\subseteq\mathbb{D}$, and $Tf=\psi\cdot(f\circ\phi)$. This completes the proof. ∎
@@ 3. Kowalski-Słodkowski’s theorem for topological spaces
where $|\alpha|=\alpha_{1}+\alpha_{2}+\dotsm+\alpha_{n}$. For convenience, we shall write $\mathcal{P}_{n}$ instead of $\mathcal{P}_{n}^{\mathcal{A}}$ if $\mathcal{A}=\mathbb{C}.$ Clearly, given any compact subset $K$ of $\mathbb{C}^{n}$, $\mathcal{P}^{\mathcal{A}}_{n}$ can be viewed as a subalgebra of the Banach algebra $C(K,\mathcal{A})$, the set of all continuous $\mathcal{A}$-valued functions with the sup norm. We say $z_{0}$ is zero of a polynomial $p\in\mathcal{P}_{n}^{\mathcal{A}}$ if $p(z_{0})=0$. For the rest of this section, let $\mathcal{A}$ represents a fixed complex Banach algebra with identity $e$.
The following result is integral to our version of the KS theorem for topological spaces (Theorem 3.2). We present it as an independent result as, apart from being interesting in its own right, it brings out the crux of the method used to prove Theorem 3.2 even more prominently. It characterizes multiplicative linear functionals on $\mathcal{P}_{n}^{\mathcal{A}}$ as non-zero complex-valued functions on $\mathcal{P}_{n}^{\mathcal{A}}$ that satisfy a condition very similar to the one assumed in the KS theorem. First, we recall the notion of a polynomial convex hull of a set in $\mathbb{C}^{n}.$ The polynomial convex hull $\widehat{D}$ of a bounded subset $D$ of $\mathbb{C}^{n}$ is consisting of all $w\in\mathbb{C}^{n}$ such that
@@
Let $\Lambda:\mathcal{P}_{n}^{\mathcal{A}}\to\mathbb{C}$ be a non-zero function with $\Lambda(0)=0$. Suppose $D$ is a bounded subset of $\mathbb{C}^{n}$ such that for every $p$ and $q$ in $\mathcal{P}^{\mathcal{A}}_{n}$ the polynomial
$\big{(}\Lambda(p)-\Lambda(q)\big{)}e-(p-q)$
has a zero in $D$. Then there exist $w\in\widehat{D}$ and a multiplicative linear functional $\phi$ on $\mathcal{A}$ such that
@@
Moreover, we assert $w\in D$ whenever
@@
For every $p,\ q\in\mathcal{P}_{n}^{\mathcal{A}}$ there exists $w_{p,q}\in D$ such that $(\Lambda(p)-\Lambda(q))e=p(w_{p,q})-q(w_{p,q})$. In particular, for each polynomial $p$, there is a point $w_{p}$ in $D$ such that $\Lambda(p)e=p(w_{p})$. This gives us
@@
Now we want to show that $\big{(}\Lambda_{1}(f)-\Lambda_{1}(g)\big{)}e-(f-g)\notin G(\overline{\mathcal{P}_{n}^{\mathcal{A}}})$ (the set of invertible elements of $\overline{\mathcal{P}_{n}^{\mathcal{A}}}$) for all $f,g\in\overline{\mathcal{P}_{n}^{\mathcal{A}}}$. Suppose there are $f$ and $g$ in $\overline{\mathcal{P}_{n}^{\mathcal{A}}}$ such that $\big{(}\Lambda_{1}(f)-\Lambda_{1}(g)\big{)}e-(f-g)\in G(\overline{\mathcal{P}_{n}^{\mathcal{A}}})$. Let $\{p_{n}\}_{n}$ and $\{q_{n}\}_{n}$ be sequences of polynomials in $\mathcal{P}_{n}^{\mathcal{A}}$ such that $p_{n}\longrightarrow f$ and $q_{n}\longrightarrow g$ in $\overline{\mathcal{P}_{n}^{\mathcal{A}}}$. Then
@@
Finally, we will establish the assertions made in the moreover part. First, suppose $n=1,$ that is, we are in the one-variable situation. Take $p(z)=ez,$ then $\Lambda(p)=w.$ Also, as we noted at the beginning of the proof, there exists a point $a\in D$ such that $\Lambda(p)e=p(a)=ea$, which implies that $w=a\in D.$
We now turn to the general $n$ case. Let $D=D_{1}\times\cdots\times D_{n}$ for some subsets $D_{1},\dots,D_{n}$ of $\mathbb{C}.$ The idea here is very similar to case $n=1$. Note that $\Lambda(r_{i})=\phi(r_{i}(w))=\phi(w_{i}e)=w_{i}$. Additionally, as mentioned earlier, some $a$ exists in $D$ such that $\Lambda(r_{i})e=r_{i}(a)=ea_{i}.$ Thus $w_{i}=a_{i}\in D_{i}$; hence, $w\in D_{1}\times\cdots\times D_{n}=D.$
@@
Also, there exists some $a\in D$ such that $\Lambda(p)e=p(a)$, which implies that $||w-b||^{2}_{\mathbb{C}^{n}}=\sum_{i=1}^{n}(\overline{w_{i}-b_{i}})(a_{i}-b_{i})=\langle{w-b,a-b}\rangle_{\mathbb{C}^{n}}$. Thus, using the Cauchy-Schwarz inequality, we infer $||w-b||_{\mathbb{C}^{n}}\leq||a-b||_{\mathbb{C}^{n}}<r.$ Thus, $w\in D$. A similar set of arguments will work when $D$ is a closed Euclidean ball in $C^{n}$. This completes the proof. ∎
@@
Let $\Lambda:\mathcal{X}\to\mathbb{C}$ be a non-zero continuous function such that $\Lambda({0})=0$ and for every $p$ and $q$ in $\mathcal{P}_{n}^{\mathcal{A}}$ the polynomial
$\big{(}\Lambda(p)-\Lambda(q)\big{)}e-(p-q)$
has a zero in $D$. Then there exist $w\in D$ and a multiplicative linear functional $\phi$ on $\mathcal{A}$ such that
@@
Clearly, the restriction $\Lambda:\mathcal{P}_{n}^{\mathcal{A}}\longrightarrow\mathbb{C}$ satisfies the hypotheses of Theorem 3.1. Then, using the moreover part of Theorem 3.1, there exist $w\in D$ and a multiplicative linear functional $\phi$ on $\mathcal{A}$ such that $\Lambda(p)=\phi(p(w))$ for every $p\in\mathcal{P}_{n}^{\mathcal{A}}$ Now the continuity of $\Lambda$ and the denseness of $\mathcal{P}_{n}^{\mathcal{A}}$ in $\mathcal{X}$ yield
@@
Remark 3.3.
The hypothesis “$\big{(}\Lambda(p)-\Lambda(q)\big{)}e-(p-q)$ has a zero in $D$” in Theorems 3.1 and 3.2 for the case when $n=1,D=\mathbb{D}\ {\rm and}\ \mathcal{A}=\mathbb{C}$ assumes nothing but that the polynomials $\big{(}\Lambda(p)-\Lambda(q)\big{)}e-(p-q)$ does not belong to the class of polynomials that are outer functions. Consequently, our hypothesis is in line with the hypothesis of the analogues of the GKZ theorem and the KS theorem recorded in the Introduction.
@@
We would like to now compare Theorem 3.2 with Theorem 1.6 ([11, Theorem 1.4]) that also is a several variables analogue of the GKZ theorem for function spaces. First, we note that Theorem 1.6 is for Banach spaces of scalar-valued functions, whereas Theorem 3.2 deals with topological spaces of vector-valued functions. Next, we compare the domains of the function spaces considered by the two theorems. Theorem 3.2, in the several variables case, is proved under the assumption that the domain is either a Euclidean ball or a Cartesian product. In contrast, it may seem that the domain of functions in Theorem 1.6 includes a big class of subsets of $\mathbb{C}^{n}$. But, we want to point out that Theorem 1.6 assumes the domain to be a “maximal domain” (hypothesis JS2 in Theorem 1.6), and if we assume the domain $D$ in Theorem 3.2 to be a maximal domain instead of being a specific set such as a Euclidean ball or a Cartesian product (as done in Theorem 3.2), then indeed with the help of Theorem 3.1, for this case also we would obtain the map $\Lambda$ to be a point evaluation for the scalar-valued case and composition of a multiplicative linear functional and a point evaluation for the vector-valued case. Moreover, Theorem 1.6 doesn’t cover Dirichlet spaces $\mathcal{D}_{\alpha}$ for $\alpha>1$ as the maximal domain for functions in these Dirichlet spaces is $\overline{\mathbb{D}^{n}}$ and not $\mathbb{D}^{n}$. But, these are covered by Theorem 3.2 as they satisfy its hypotheses.
@@
For $n,m\in\mathbb{N}$ and $0<p\leq\infty$, let $\Lambda:H^{p}(\mathbb{D}^{n},\mathbb{C}^{m})\to\mathbb{C}$ be a non-zero continuous (weak-star continuous for $p=\infty$) function such that $\Lambda({0})=0$ and for polynomials $p$ and $q$ in $\mathcal{P}^{\mathbb{C}^{m}}_{n}$ the polynomial $\big{(}\Lambda(p)-\Lambda(q)\big{)}{\bf 1}-(p-q)$ has a zero in $\mathbb{D}^{n}$. Then there exist $w\in\mathbb{D}^{n}$ and $1\leq k\leq m$ such that
@@
For $0<p,q\leq\infty,$ let $T:H^{p}(\mathbb{D}^{n})\to H^{q}(\mathbb{D}^{m})$ be a continuous function, where we consider weak-start topology on $H^{p}(\mathbb{D}^{n})$ when $p=\infty.$ Suppose $T(0)=0,\ T(1)(w)\neq 0$ for each $w\in\mathbb{D}^{m}$. If given any pair $p,q$ of complex-valued polynomials in $n$-variables and $w\in\mathbb{D}^{m}$ the polynomial
$\big{(}Tp(w)-Tq(w)\big{)}1-\big{(}T1(w)\big{)}(p-q)$
has a zero in $\mathbb{D}^{n}$, then there exist $\psi\in H^{q}(\mathbb{D}^{m})$ and an analytic function $\phi:\mathbb{D}^{m}\to\mathbb{D}^{n}$ such that